OpenStudy (anonymous):

Evaluate the expression without using a calculator. 71 • 9 A. 549 B. 639 C. 719 D. 720

OpenStudy (cwrw238):

first multiply 1 by 9 then multiply 70 by 9 and add the 2 results

OpenStudy (anonymous):

so ummm what is the answer

OpenStudy (anonymous):

the answer is 630 right :D

OpenStudy (anonymous):

71*9=(70+1)*9=70*9+1*9=630+9=?

OpenStudy (cwrw238):

well 1*9 = 9 and 70 * 9 = 630

OpenStudy (anonymous):

639???????

random231 (random231):

yeah 639

OpenStudy (anonymous):

360+9=639

OpenStudy (anonymous):

Yes it's "b" 639

OpenStudy (anonymous):

What is the solution to the equation? 7(5 + 9) = 7 · x + 7 · 9 A. x = 1 B. x = 5 C. x = 7 D. x = 9

OpenStudy (anonymous):

a(b+c)=ab+ac